P,Q,R is a point on the side of the triangle AB, BC and CA respectively. if P, Q and R divides the sides AB, BC and CA in the ratio m:1. Find the centroid of the triangle ABC.

Grade:11

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Chanchal ,

let A = (a1 ,a2) ; B =(b1, b2) ; C =(c1, c2)

p = (mb1+a1/m+1 , mb2+a2/m+1)  , Q = (mc1+b1/m+1 , mc2+b2/m+1) , R = (ma1+c1/m+1 , ma2+c2/m+1)

centroid  = p+Q+R/3

   = (m(a1+b1+c1)+(a1+b1+c1)/(m+1)) /3, m(a2+b2+c2)+(a2+b2+c2)/(m+1)/3

   = (a1+b1+c1)/3 ,(a2+b2+c2)/3

i.e same as the triangle ABC .

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum..